OpenStudy (anonymous):

8^(sqrt(x)) = 4^(sqrt(x)) + 1 find x

OpenStudy (anonymous):

\[8^{\sqrt{x}}=4^{\sqrt{x}}+1\] Let's find \[y=2^\sqrt{x}\] first \[y^3=y^2+1 \rightarrow y^3-y^2-1=0\]

OpenStudy (jamesj):

Let u = 2^sqrt(x). Then notice that u^2 = 4^sqrt(x) and u^3 = 8^sqrt(x) So u^3 = u^2 + 1 .... u^3 - u^2 - 1 = 0

OpenStudy (jamesj):

The unfortunate thing now is that this equation looks simple but doesn't have a particularly nice solution.

OpenStudy (anonymous):

u^2 = 2^(2(sqrt(x))